### v1

# definitionDefinition[section]

Abstract
In (Dzanic, J. Comp. Phys., 508:113010, 2024), a limiting approach for high-order discontinuous Galerkin schemes was introduced which allowed for imposing constraints on the solution continuously (i.e., everywhere within the element). While exact for linear constraint functionals, this approach only imposed a sufficient (but not the minimum necessary) amount of limiting for nonlinear constraint functionals. This short note shows how this limiting approach can be extended to allow exactness for general nonlinear quasiconcave constraint functionals through a nonlinear limiting procedure, reducing unnecessary numerical dissipation. Some examples are shown for nonlinear pressure and entropy constraints in the compressible gas dynamics equations, where both analytic and iterative approaches are used.
keywords:
[…]
1 Introduction
The use of high-order discontinuous Galerkin (DG) schemes for simulating transport-dominated physics often requires additional numerical stabilization to ensure the discrete solution is structure-preserving (i.e., is well-behaved and satisfies physical constraints). One such way of achieving this desirable behavior is applying some form of a posteriori limiting on the solution, where the high-order DG approximation is blended with a secondary, more robust (but generally less accurate) approximation to yield a bounds-preserving (i.e., constraint-satisfying) solution. This form of limiting has been widely used for simulating complex physical systems, particularly ones of predominantly hyperbolic nature, where numerical solutions must abide by known physical constraints (e.g., positivity of density and pressure in gas dynamics). However, a potential drawback of these limiting methods is that they are typically performed at discrete nodal locations for the DG solution, which only ensures that the given constraints are satisfied at those points. While adequate for many problems, the use of discretely bounds-preserving limiting may not be robust enough for applications where the DG solution must be evaluated at arbitrary locations (e.g., coupled meshes/solvers, remapping in arbitrary Lagrangian-Eulerian methods, adaptive mesh refinement, etc.). In such scenarios, the limited solution may still violate the constraints at these arbitrary points, causing the failure of the numerical scheme.
A potential remedy for this problem was introduced in Dzanic [1] through a limiting approach that ensures the limited solution is continuously bounds-preserving (i.e., across the entire solution polynomial). The approach relied on a novel functional for computing the limiting factor for the “squeeze” limiter of Zhang and Shu [2], such that if the minimum of this limiting functional within an element was found (via a spatial optimization algorithm), the limited solution was guaranteed to be continuously bounds-preserving for any arbitrary quasiconcave constraint functional. The underlying mathematical basis of this approach was a linearization of the constraint functionals being enforced on the solution. For linear constraints (e.g., maximum principle on scalar solutions, positivity of density in gas dynamics, etc.), the computed limiting factor was “exact” (i.e., it was the minimum necessary amount of limiting). However, for nonlinear constraint functionals, the linearization could only ensure that the computed limiting factor was sufficient, such that a smaller limiting factor (i.e., less numerical dissipation) could still yield a solution which satisfied the constraints continuously. Therefore, the limiting approach was suboptimal for nonlinear constraints commonly encountered in hyperbolic systems of equations as it was not applying only the minimum amount of limiting necessary.
The purpose of this short note is to show how this limiting functional can be further modified such that the computed limiting factor is exact for arbitrary nonlinear quasiconcave functionals, reducing unnecessary numerical dissipation. While this proposed modification can be applied to general hyperbolic conservation laws, we focus specifically on the compressible Euler equations which admit higher-order and nonlinear constraints for pressure and entropy. Some examples are shown for the new approach, where both analytic and iterative approaches can be applied.
2 Preliminaries
In this section, some preliminaries are presented for the proposed modification, which briefly summarize the limiting approach introduced in Dzanic [1].
2.1 Governing equations and constraints
The exemplar hyperbolic conservation law used for this work is the compressible Euler equations for gas dynamics in $d$ dimensions, written in conservation form as
$\partial_{t}\mathbf{u}(\mathbf{x},t)+\boldsymbol{\nabla}\cdot\mathbf{F}(\mathbf{u})=0,$ (1)
where
$\mathbf{u}=\begin{bmatrix}\rho\\ \mathbf{m}\\ E\end{bmatrix}\quad\mathrm{and}\quad\mathbf{F}(\mathbf{u})=\begin{bmatrix}\mathbf{m}^{T}\\ \mathbf{m}\otimes\mathbf{v}+P\mathbf{I}\\ (E+P)\mathbf{v}^{T}\end{bmatrix}.$ (2)
Here, $\rho$ is the density, $\mathbf{m}$ is the momentum, and $E$ is the total energy. The symbol $\mathbf{I}$ denotes the identity matrix in $\mathbb{R}^{d\times d}$, $\mathbf{v}=\mathbf{m}/\rho$ denotes the velocity, and $P$ denotes the pressure, computed as
$P=(\gamma-1)\rho e=(\gamma-1)\left(E-\frac{1}{2}\mathbf{m}{\cdot}\mathbf{m}/\rho\right),$ (3)
where $\gamma=1.4$ is the specific heat ratio and $e$ is the specific internal energy.
The solution of the Euler equations is endowed with a convex invariant set corresponding to the positivity of density ($\rho\geq 0$), positivity of pressure/internal energy ($P,\rho e\geq 0$), and a minimum principle on the specific physical entropy ($\sigma\geq\sigma_{\min}$), where $\sigma=P\rho^{-\gamma}$ [3]. This can be represented by the positivity of a set of constraint functionals as
$g_{1}(\mathbf{u})=\rho-\rho_{\min},\quad g_{2}(\mathbf{u})=P-P_{\min},\quad\text{and}\quad g_{3}(\mathbf{u})=\sigma-\sigma_{\min},$ (4)
where small tolerances ($\rho_{\min}$, $P_{\min}$) are included for numerical stability purposes. This set of constraints is particularly interesting as they each have a unique mathematical character, consisting of a linear density constraint, a quadratic pressure/internal energy constraint, and a nonlinear entropy constraint. In many limiting approaches, the entropy constraint is neglected and only positivity of density/pressure is enforced (e.g., positivity-preserving limiters).
2.2 Discontinuous Galerkin methods
The underlying numerical method for this approach is the discontinuous Galerkin scheme [4], where the solution $\mathbf{u}(\mathbf{x})$ within each element of the mesh is represented by a set of $n_{s}$ basis functions as
$\mathbf{u}(\mathbf{x})=\sum_{i=1}^{n_{s}}\mathbf{u}_{i}\phi_{i}(\mathbf{x})\subset V_{h},$ (5)
where $\phi_{i}(\mathbf{x})$ are the basis functions, $\mathbf{u}_{i}$ are their associated coefficients, and $V_{h}$ is the piece-wise polynomial space spanned by the basis functions. We consider a semi-discrete weak formulation of a hyperbolic conservation law in the form of
$\sum_{k=1}^{N}\left\{\int_{\Omega_{k}}\partial_{t}\mathbf{u}{\cdot}\mathbf{w}\ \mathrm{d}V+\int_{\partial\Omega_{k}}\hat{\mathbf{F}}(\mathbf{u}^{-},\mathbf{u}^{+},\mathbf{n}){\cdot}\mathbf{w}\ \mathrm{d}S-\int_{\Omega_{k}}\mathbf{F}(\mathbf{u}){\cdot}\nabla\mathbf{w}\ \mathrm{d}V\right\}=0,$ (6)
where $\mathbf{w}(\mathbf{x})\subset V_{h}$ is a test function and $\hat{\mathbf{F}}(\mathbf{u}^{-},\mathbf{u}^{+},\mathbf{n})$ is a numerical interface flux. Furthermore, we define the element-wise mean for an arbitrary element $\Omega_{k}$ as
$\overline{\mathbf{u}}=\frac{\int_{\Omega_{k}}\mathbf{u}(\mathbf{x})\ \mathrm{d}\mathbf{x}}{\int_{\Omega_{k}}\mathrm{d}\mathbf{x}},$ (7)
which, under some relatively minor assumptions on the numerical scheme, generally preserves convex invariants of hyperbolic systems, such that it may be used as secondary bounds-preserving approximation for a limiting approach (see Zhang and Shu [2], Zhang and Shu [5], and derived works).
2.3 Continuously bounds-preserving limiting
The proposed limiting approach in Dzanic [1] relies on the a posteriori “squeeze” limiter of Zhang and Shu [2], which linearly contracts the high-order DG solution towards the element-wise mean based on a limiting factor $\alpha$ as
$\widehat{\mathbf{u}}(\mathbf{x})=(1-\alpha)\mathbf{u}(\mathbf{x})+\alpha\overline{\mathbf{u}}=\mathbf{u}(\mathbf{x})+\alpha\left(\overline{\mathbf{u}}-\mathbf{u}(\mathbf{x})\right).$ (8)
The goal of the proposed approach is to find a limiting factor such that
$g(\widehat{\mathbf{u}}(\mathbf{x}))\geq 0\ \forall\ \mathbf{x},$ (9)
for any arbitrary quasiconcave constraint functional $g(\mathbf{u})$. This approach differed from standard a posteriori limiting methods for DG schemes in that it enforced constraints continuously instead of at discrete nodal locations (here, $\forall\ \mathbf{x}$ is used to refer to all locations within the element). It was shown in Dzanic [1] that if one introduces the modified limiting functional $h(\mathbf{u})$ as
$h(\mathbf{u})=\begin{cases}h^{+}(\mathbf{u}),\quad\text{if }g(\mathbf{u})\geq 0,\\ h^{-}(\mathbf{u}),\quad\text{else},\end{cases}$ (10)
where
$h^{+}(\mathbf{u})=\frac{g(\mathbf{u})}{g(\overline{\mathbf{u}})}\quad\text{and}\quad h^{-}(\mathbf{u})=\frac{g(\mathbf{u})}{g(\overline{\mathbf{u}})-g(\mathbf{u})},$ (11)
then setting the limiting factor $\alpha$ as
$\alpha=\max\left[0,-\underset{\mathbf{x}}{\min}\ h(\mathbf{u})\right],$ (12)
guarantees a limiting scheme that is continuously bounds-preserving.
In this approach, one can define “exactness” in the limiting as finding the minimum necessary value of $\alpha$ such that Eq. 9 is satisfied, which, for high-order solutions that initially violate the constraints, can be expressed as $\underset{\mathbf{x}}{\min}\ g(\widehat{\mathbf{u}}(\mathbf{x}))=0$. However, for nonlinear quasiconcave constraint functionals, the linearization used in the definition of $h^{-}(\mathbf{u})$ can only ensure that the limiting is sufficient, i.e., $\underset{\mathbf{x}}{\min}\ g(\widehat{\mathbf{u}}(\mathbf{x}))\geq 0$.
3 Proposed modifications
The primary motivation of this work is to introduce a modification to the definition of $h^{-}(\mathbf{u})$ such that exactness can also be ensured for nonlinear functionals, reducing unnecessary numerical dissipation in the limiting procedure. Exactness for the limiting factor can be represented as finding the minimum necessary value of $\alpha$ such that constraints are satisfied, i.e.,
$\underset{\alpha\geq 0}{\arg\min}\ g\left((1-\alpha)\mathbf{u}(\mathbf{x})+\alpha\overline{\mathbf{u}}\right)\geq 0\ \forall\ \mathbf{x},$
which is equivalent to finding the spatial maximum of the necessary limiting factor at every point in the element, i.e.,
$\underset{\mathbf{x}}{\max}\ \left[0,\alpha^{*}(\mathbf{x})\right],\quad\text{where}\quad g\left((1-\alpha^{*}(\mathbf{x}))\mathbf{u}(\mathbf{x})+\alpha^{*}(\mathbf{x})\overline{\mathbf{u}}\right)=0.$
To this end, we propose achieving this by replacing the linearized formulation of $h^{-}(\mathbf{u})$ is Eq. 10, denoted by $h^{-}_{L}(\mathbf{u})$, with a better (nonlinear) approximation of the necessary limiting factor, denoted by $h^{-}_{NL}(\mathbf{u})$, which mimics the above condition as
$h_{NL}^{-}(\mathbf{u}(\mathbf{x}))=-\alpha^{*}(\mathbf{x}):g(\widehat{\mathbf{u}}(\mathbf{x}))=g\left((1-\alpha^{*}(\mathbf{x}))\mathbf{u}(\mathbf{x})+\alpha^{*}(\mathbf{x})\overline{\mathbf{u}}\right)=0$ (13)
It can then be trivially shown following the proof in Dzanic [1] that setting the limiting factor as Eq. 12 using this modified formulation ensures $\underset{\mathbf{x}}{\min}\ g(\widehat{\mathbf{u}}(\mathbf{x}))=0$ if the high-order solution is bounds-violating and $g(\widehat{\mathbf{u}}(\mathbf{x}))=g(\mathbf{u})$ if the high-order solution is bounds-preserving, achieving exactness for arbitrary nonlinear constraint functionals. Furthermore, it also ensures $C^{0}$ continuity in $h(\mathbf{u})$ as $h_{NL}^{-}(\mathbf{u})=h^{+}(\mathbf{u})$ when $g(\mathbf{u})=0$.
However, this formulation requires solving an intersection problem for the zero contour of the constraint functional. For linear constraints, it can be seen that this reduces to the linearized formulation, i.e., $h^{-}_{NL}(\mathbf{u})=h^{-}_{L}(\mathbf{u})$. For quadratic constraints such as pressure, this higher-order limiting approach also admits an analytic solution, where $\alpha^{*}$ can be computed as the (positive) root of the quadratic equation
$A\left(\alpha^{*}\right)^{2}+B\alpha^{*}+C=0,$
with the coefficients taking on the values
$A=\Delta\rho\Delta E-\frac{1}{2}\Delta\mathbf{m}\cdot\Delta\mathbf{m},\quad B=E\Delta\rho+\rho\Delta E-\mathbf{m}\cdot\Delta\mathbf{m}-\Delta\rho\frac{P_{\min}}{\gamma-1},\quad C=\rho E-\frac{1}{2}\mathbf{m}\cdot\mathbf{m}-\rho\frac{P_{\min}}{\gamma-1},$
for positive pressure constraints, where $\Delta\mathbf{u}=\overline{\mathbf{u}}-\mathbf{u}$. Similar closed-form expressions can be used for higher-order constraints (e.g., cubic) which admit analytic solutions to the intersection problem in Eq. 13.
If one also wants to ensure exactness for nonlinear constraints such as entropy which do not have an analytic solution to the intersection problem, it can be straightforwardly solved using simple iterative root-bracketing approaches such as the bisection or Illinois method, which is similar to discrete limiting approaches with entropy-based constraints [6]. The quasiconcavity of the constraint functionals ensures that this root-bracketing problem is well-behaved, i.e., the is solution is bounded by $0\leq\alpha^{*}\leq 1$ and is unique. Furthermore, one can apply a stricter upper bound for the initial bracket by using the linearized formulation in Eq. 11, further reducing the computational cost.
4 Results
The modified functional was implemented within the same numerical framework as in Dzanic [1], using identical parameters and optimization approaches (for more information, the reader is referred to Section 3 of Dzanic [1]). The intersection problem for entropy constraints was solved using 5 iterations of the Illinois method. Minimum density and pressure values were set as $\rho_{\min}=P_{\min}=10^{-11}$.
To highlight the improvements of the proposed modification, we consider first a pathological example of limiting a static discontinuity within one element, where a discontinuity is placed in the center of the element with the left/right states set as
$\left[\rho,u,P\right]^{T}=\begin{cases}\left[1,1,2P_{\min}\right]^{T},&\mbox{if }x\leq 0.5,\\ \left[3,3,1\right]^{T},&\mbox{else}.\end{cases}$ (14)
A 9th-order polynomial DG approximation was initialized by interpolating the solution on the Gauss–Lobatto nodes spaced along the domain $\Omega=[0,1]$, and constraints were enforced for positivity of density/pressure as well as a minimum entropy principle. An arbitrary minimum entropy value was set as $\sigma_{\min}=0.1$. These conditions yield various properties for the unlimited solution to showcase the behavior of the new approach, with a continuously bounds-preserving density field, a discretely (but not continuously) bounds-preserving pressure field, and a discretely bounds-violating entropy field.
[…]
(a) Pressure
[…]
(b) Entropy
Figure 1: Comparison of the linear and nonlinear limiting approaches for pressure (left) and entropy (right) constraints for a static discontinuity problem computed with a $\mathbb{P}_{9}$ approximation. Pressure limiting computed using analytic formulation, entropy limiting computed using iterative formulation. Dotted line represents zero contour for the constraint functional.
As the density field was already continuously bounds-preserving, the limiter was first applied to the pressure field. A comparison of the linearized formulation and the proposed nonlinear formulation is shown in Fig. 1. It can be seen that the proposed nonlinear formulation drastically reduces the numerical dissipation in the scheme, with the limited solution significantly closer to the unlimited solution in comparison to the linearized formulation. Furthermore, it can be seen that the new limiting approach is exact in the sense that the minimum pressure of the limited solution was on the order of $P_{\min}$. Similar results could be seen for the entropy constraints, also shown in Fig. 1. Due to the stronger nonlinearity of the constraint functional, the linearized approach showed even more unnecessary dissipation than with the pressure constraint, whereas the proposed nonlinear approach was exact and showed a significant reduction in the numerical dissipation.
A more quantitative evaluation of the proposed modification was performed with a near-vacuum isentropic Euler vortex which possesses an analytic solution for comparison. The problem consists of a smooth vortex on a periodic domain $\Omega=[-10,10]^{2}$, with the initial solution set as
$\begin{bmatrix}\rho\\ u\\ v\\ P\end{bmatrix}=\begin{bmatrix}P^{\frac{1}{\gamma}}\\ \frac{S}{2\pi R}(y-y_{0})\phi(\mathbf{x})\\ 1-\frac{S}{2\pi R}(x-x_{0})\phi(\mathbf{x})\\ \frac{1}{\gamma M^{2}}\left(1-\frac{S^{2}M^{2}(\gamma-1)}{8\pi^{2}}\phi(\mathbf{x})^{2}\right)^{\frac{\gamma}{\gamma-1}}\end{bmatrix},\quad\mathrm{where}\quad\phi(\mathbf{x})=\exp{\left(\frac{1-\|\mathbf{x}-\mathbf{x}_{0}\|_{2}^{2}}{2R^{2}}\right)}.$
The parameters are set to $R=1.5$, $M=0.4$, and $S=28.11711$, which yield a near-vacuum state at the vortex peak with a minimum density of $\rho=8{\cdot}10^{-9}$ and minimum pressure of $P=2{\cdot}10^{-11}=2P_{\min}$. One flow-through of the domain ($t=20$) was computed using both a $\mathbb{P}_{4}$ and $\mathbb{P}_{5}$ approximation using meshes of varying resolution with only positivity-preserving constraints, after which the $L^{\infty}$ norm of the pressure error (computed at solution nodes) was compared from the solutions obtained using the original linear limiting approach and the proposed nonlinear limiting approach. The comparisons for both the $\mathbb{P}_{4}$ and $\mathbb{P}_{5}$ approximations are shown in Section 4 and Section 4, respectively. It can be seen that the proposed modification significantly reduces the overall error in the pressure field, with both approximation orders showing decreases between $20-70\%$ across the varying levels of mesh resolution. These results highlight the reduced numerical dissipation that stems from the modified constraint functional proposed in this note.
Method $N=20^{2}$ $N=30^{2}$ $N=40^{2}$ $N=50^{2}$ $N=60^{2}$ $N=70^{2}$
Linear limiting $2.86\times 10^{-1}$ $2.93\times 10^{-2}$ $1.61\times 10^{-2}$ $7.17\times 10^{-3}$ $3.40\times 10^{-3}$ $2.34\times 10^{-3}$
Nonlinear limiting $1.55\times 10^{-1}$ $2.36\times 10^{-2}$ $9.88\times 10^{-3}$ $4.33\times 10^{-3}$ $1.61\times 10^{-3}$ $6.27\times 10^{-4}$
Error reduction -45.9% -19.4% -38.5% -39.6% -52.8% -73.2%
[…]
table Comparison of the $L^{\infty}$ norm of the pressure error using linear and nonlinear limiting for the near-vacuum isentropic Euler vortex after one flow-through of the domain ($t=20$) computed with a $\mathbb{P}_{4}$ approximation.
Method $N=20^{2}$ $N=30^{2}$ $N=40^{2}$ $N=50^{2}$ $N=60^{2}$ $N=70^{2}$
Linear limiting $1.15\times 10^{-1}$ $2.02\times 10^{-2}$ $3.58\times 10^{-3}$ $3.89\times 10^{-3}$ $1.97\times 10^{-3}$ $4.18\times 10^{-4}$
Nonlinear limiting $4.97\times 10^{-2}$ $1.66\times 10^{-2}$ $2.78\times 10^{-3}$ $1.24\times 10^{-3}$ $5.36\times 10^{-4}$ $2.43\times 10^{-4}$
Error reduction -56.5% -21.4% -27.1% -70.1% -72.0% -37.7%
[…]
table Comparison of the $L^{\infty}$ norm of the pressure error using linear and nonlinear limiting for the near-vacuum isentropic Euler vortex after one flow-through of the domain ($t=20$) computed with a $\mathbb{P}_{5}$ approximation.
5 Concluding remarks
We proposed an improvement to the continuously bounds-preserving limiting approach presented in Dzanic [1], which allows for the approach to achieve the exact amount of limiting necessary for arbitrary nonlinear quasiconcave constraint functionals as opposed to just a sufficient amount. The modification relies on replacing the linearization in limiting functional with an intersection/root-finding problem, which may be computed analytically for some constraints and numerically for others. Some examples were shown for the compressible Euler equations, showing the reduced numerical dissipation and increased accuracy of the proposed approach.
Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE–AC52–07NA27344 and the LLNL-LDRD Program under Project tracking No. 24–ERD–050. Release number LLNL–JRNL–863292–DRAFT.